OpenStudy (loser66):

What's wrong with this logic? t1 = -25\(^0C\) t2 = 40 \(^0C\) \(\triangle t = t2-t1= 65^0C = 338.15 K\) but if I convert the t1, t2 at the beginning, \(\triangle t = 65 K\) Please, explain m

OpenStudy (loser66):

@Mertsj

OpenStudy (anonymous):

You want to know why the change in temperature works either way?

OpenStudy (anonymous):

Oh I see the issue here.

OpenStudy (loser66):

I want to know why the results are different

OpenStudy (mertsj):

She wants to know why if you convert to Kelvin before finding delta t, you get delta t = 65 degrees kelvin. If you convert to kelvin after finding delta t, you get 338 degrees kelvin

OpenStudy (mertsj):

\[(40+273)-(-25+273)=65 degrees Kelvin\] \[40-(-25)=65 ^{o}C=65+273=338^{o}K\]

OpenStudy (loser66):

Thank you @Mertsj That's what I mean

OpenStudy (anonymous):

Consider \(f(x) = 4x\) and \(g(x) = 4x+10\). Notice \(g(x) = f(x) + 10\) \(\Delta f= 4(a-b)\) while \(\Delta g= 4(a-b)\). If you try convert after finding the difference, you would get \(\Delta f =4(a-b) \implies \Delta g =\Delta f + 10 = 4(a-b)+10 \)

OpenStudy (anonymous):

So basically, interval lengths don't convert.

OpenStudy (loser66):

I am with you at the last comment only. :) However, I got it.

OpenStudy (loser66):

Thank you very much @wio

OpenStudy (anonymous):

No problem I guess? Actually, here is how look at it: The function \(f(x) = x+273\) converts C to K. You are esentially saying: \[ f(x_2)-f(x_1) = f(x_2-x_1) \] In reality \[ f(x_2)-f(x_1) = x_2+273 -(x_1+273) = x_2-x_1 \]and \[ f(x_2-x_1) = (x_2-x_1) + 273 \]

OpenStudy (loser66):

Yes, I got it, the gap between the two are the same no matter what the unit is, right? in this case 65 units in difference for both them (something like the situation of 3hours and 3 o'clock) :)

OpenStudy (anonymous):

Yeah, I'm just trying to pinpoint the exact error and in general it is \(\Delta f = f(\Delta x)\) or something.

OpenStudy (loser66):

I have another question about degenerate and non-degenerate, can you help me?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

what's the context of degenerate?

OpenStudy (loser66):

what do they exactly mean? not from definition

OpenStudy (anonymous):

What context?

OpenStudy (loser66):



OpenStudy (loser66):

In linear functional, generate can be expressed as span. Can I expand this logic to bilinear form for degenerate ? something likes span the tensor product set?

OpenStudy (anonymous):

So \[ f(x,y) = (4-y)g(x,y) \]would likely be a degenerate form since \(y=4\) would probably zero it out?

OpenStudy (loser66):

in which field?

OpenStudy (anonymous):

I'm talking about your definition.

OpenStudy (loser66):

oh the definition from the book: A bilinear form w on U\(\otimes\)V is degenerate if, as a function of one of its two arguments, it vanishes identically for some non-zero value of its other argument; otherwise it is non-degenerate.

OpenStudy (anonymous):

I am not sure. I'm not completely familiar.